OpenStudy (anonymous):

Write an equation in​ slope-intercept form of the line that satisfies the given conditions. Through ​(-6−6​); parallel to -x+6y=36

OpenStudy (alexandervonhumboldt2):

first lets convert this equation to slope intercept form. do you know where to start?

pooja195 (pooja195):

Ok y=mx+b m=slope b=y intercept

OpenStudy (anonymous):

y=mx+b y=m(-6)-6

OpenStudy (anonymous):

@TheSmartOne

pooja195 (pooja195):

Wait lets do this a different way.

pooja195 (pooja195):

Can you isolate y? -x+6y=36

pooja195 (pooja195):

do you know how?

OpenStudy (anonymous):

do i put the (-6,-6) into the equation -x+6y=36

pooja195 (pooja195):

No ignore that

OpenStudy (anonymous):

-1(-6)+6(-6)=36

pooja195 (pooja195):

Nope We need to make an equation -x+6y=36 We start by adding x to both sides (since we want to isolate x) -x+x+6y=36+x This leaves us with 6y=36+x we need to divide all of this by 6 so we isolate y 6y/6=36/6+x what do we have left?

OpenStudy (anonymous):

y=6+x

pooja195 (pooja195):

x/6 would just be 1/6 since x is just 1x y=6+1/6x can you put that in slope intercept form? y=mx+b m=slope b=y intecet

OpenStudy (anonymous):

y=1/6x+6

OpenStudy (anonymous):

is that right @pooja195

pooja195 (pooja195):

Yes but we are not done yet any ideas on what the next step is? :)

OpenStudy (anonymous):

no idea

pooja195 (pooja195):

so the lines are parallel means same slope we just steal the slope form y = 1/6x + 6 and we get y = 1/6x + b now we plug in (-6,-6) and solve for b and whatever you get for b you plug into y = 1/6x + b and you're done

TheSmartOne (thesmartone):

steal the slope from*

pooja195 (pooja195):

Do you understand? @kirstenc73093

OpenStudy (anonymous):

y=-7

pooja195 (pooja195):

Dont over think this! y = 1/6x + b B=y intercept in the given coordinates (-6,-6) Which is the y value?

OpenStudy (anonymous):

-6 is the y

pooja195 (pooja195):

good so plug that into the equation

TheSmartOne (thesmartone):

don't forget to plug in the x-value too!

OpenStudy (anonymous):

y=1/6(-6)+-6

TheSmartOne (thesmartone):

no, like this: -6 = 1/6 (-6) + b solve for b :)

TheSmartOne (thesmartone):

\(\sf\Large -6 = \frac{1}{6} (-6) + b\)

OpenStudy (anonymous):

b=-5

pooja195 (pooja195):

Good now make that into slope intercept from m=1/6 y intercept= -5

OpenStudy (anonymous):

-5=1/6+5

pooja195 (pooja195):

o.O slope intercept form y=mx+b m=slope b=y intercept

OpenStudy (anonymous):

y=1/6x-5

pooja195 (pooja195):

Good job! :)

OpenStudy (anonymous):

is that the end?

pooja195 (pooja195):

yes

TheSmartOne (thesmartone):

yay! Great job @pooja195 and @kirstenc73093 :D

OpenStudy (anonymous):

thanks @TheSmartOne and @pooja195

TheSmartOne (thesmartone):

Don't forget to rate the QH! :D

OpenStudy (anonymous):

oh don't worry I did! :)

pooja195 (pooja195):

thanks! :)

TheSmartOne (thesmartone):

:)